# Quantum Algorithms in a Superposition of Spacetimes

### v1

@@ 7 $\mathsf{GapCVP}_{O\left(n^{2}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
$\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}\right)\right\rVert_{TV}\leq n\cdot\left\lVert\mathcal{D}\left(\mathbf{e}_{1}\right)-\mathcal{D}\left(D\left(\tilde{B}\right)\right)\right\rVert_{TV}\enspace,$
where $\mathbf{e}_{1}$ is the distribution of the first coordinate of the vector $\mathbf{e}$ (which distributes identically for all of its coordinates $i\in[n]$). Recall that $\mathbf{e}_{1}$ is a random variable that’s the sum of $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-2^{\beta},2^{\beta}-1\right)}$ and $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$. By Lemma 7.1, there exists a positive absolute constant $c_{0}\in\mathbb{N}$ such that $\left\lVert\mathcal{D}\left(\mathbf{e}_{1}\right)-\mathcal{D}\left(D\left(\tilde{B}\right)\right)\right\rVert_{TV}\leq c_{0}\cdot\frac{1}{\sqrt{\kappa}}$.
@@
$\left\lVert\left(S\right)-\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace,$
@@
  • The expectation of $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$ is $\mu_{i,0}=-\frac{1}{2}$ and the expectation of $\mathcal{U}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is $\mu_{i,1}=\frac{1}{2}$.
  • The variance of both $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$, $\mathcal{U}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is the same: $\sigma_{i,b}^{2}=\frac{\left(2^{\beta+1}+1\right)^{2}-1}{12}=\frac{2^{2\beta+2}+2^{\beta+2}}{12}=\Theta\left(2^{2\beta}\right)$.
@@
  • The third moment of $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$, $\mathcal{U}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is of the same order:
    $\gamma_{i,0}=\gamma_{i,1}:=\frac{\mathbb{E}_{\mathcal{U}_{\left(-2^{\beta},2^{\beta}-1\right)}}\left(|\mathcal{U}_{\left(-2^{\beta},2^{\beta}\right)}-\mu_{i}|^{3}\right)}{\sigma^{3}}=\Theta\left(\frac{1}{\kappa\sqrt{\kappa}}\right)\enspace.$
@@
  • For every $i\in[\kappa/2],b\in\{0,1\}$, the random variable $S^{(i,b)}:=S-\mathcal{U}^{\left(i,b\right)}$ is the the discrete uniform sum of $\kappa-1$ uniform distributions, instead of summing all $\kappa$ samples as in $S$.
@@
$\left\lVert\left(S\right)-\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)\right\rVert_{TV}$
$\leq\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\gamma_{i,b}+\frac{2\cdot\sigma_{i,b}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\left(S^{(i,b)}\right)-\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
@@
where $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$. We analyze each of the three summands separately, from last to first. As for the third (and last) summand, $\frac{1}{\sigma\cdot 2\sqrt{2\pi}}\leq O\left(\frac{1}{\sqrt{\kappa}\cdot 2^{\Omega\left(\beta\right)}}\right)$ because $\sigma\leq\Omega\left(\sqrt{\kappa}\cdot 2^{\Omega\left(\beta\right)}\right)$. The second summand is $\left(5+3\sqrt{\frac{\pi}{8}}\right)\cdot\gamma\leq O\left(\frac{1}{\sqrt{\kappa}}\right)$ because $\gamma\leq O\left(\frac{1}{\sqrt{\kappa}}\right)$. Overall the sum of the third and second summands together is $O\left(\frac{1}{\sqrt{\kappa}}\right)$.
@@
$\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\gamma_{i,b}+\frac{2\cdot\sigma_{i,b}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\left(S^{(i,b)}\right)-\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
$=\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{i,b}+\frac{\sigma_{i,b}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\left(S^{(i,b)}\right)-\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
=(*$\underset{(*)}{=}\kappa\cdot\Theta\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{1,0}+\frac{\sigma_{1,0}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\left(S^{(1,0)}\right)-\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$\leq O\left(2^{\beta}\cdot\frac{1}{\sqrt{\kappa}}\cdot\left\lVert\left(S^{(1,0)}\right)-\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)+O\left(\frac{1}{\kappa}\cdot\left\lVert\left(S^{(1,0)}\right)-\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$\leq O\left(2^{\beta}\cdot\frac{1}{\sqrt{\kappa}}\cdot\left\lVert\left(S^{(1,0)}\right)-\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)\enspace,$
where the equality (*)(*)( * ) follows because for every $i\in[\kappa/2],b\in\{0,1\}$, $\gamma_{i,b}=\Theta\left(\gamma_{1,0}\right)$, $\sigma_{i,b}=\sigma_{1,0}$, and
$\left\lVert\left(S^{(i,b)}\right)-\left(S^{(i,b)}+1\right)\right\rVert_{TV}=\Theta\left(\left\lVert\left(S^{(1,0)}\right)-\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)\enspace.$
Finally, by Fact 7.1, the above is bounded by $O\left(\frac{1}{\sqrt{\kappa}}\right)$.
@@
$\left\lVert\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\left(D_{\tilde{B}}\right)\right\rVert_{TV}$
$\leq\left\lVert\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\tilde{B}\right)\right)-\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)-\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace.$
@@
Let $S$ and $S^{\left(i,b\right)}$ as defined in the statement of Lemma 7.1. Then,
$\left\lVert\left(S^{\left(i,b\right)}\right)-\left(S^{\left(i,b\right)}+1\right)\right\rVert_{TV}\leq O\left(2^{-\beta}\right)\enspace.$
@@
$=e^{-\frac{\pi\cdot\left(\left\lVert\mathbf{e}_{0}\right\rVert^{2}+2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}\cdot B\right)}{2\cdot B^{2}}}\cdot\sum_{\mathbf{e}\in\mathbb{Z}^{n}_{\left(-B,\cdots,-1,0,1,\cdots,B\right)}}\frac{1}{P}\cdot e^{-\frac{\pi 2\cdot\left\lVert\mathbf{e}\right\rVert^{2}}{2\cdot B^{2}}}$

### v2

@@ 7 $\mathsf{GapCVP}_{O\left(n^{2}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
$\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}\right)\right\rVert_{TV}\leq n\cdot\left\lVert\mathcal{D}\left(\mathbf{e}_{1}\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\enspace,$
where $\mathbf{e}_{1}$ is the distribution of the first coordinate of the vector $\mathbf{e}$ (which distributes identically for all of its coordinates $i\in[n]$). Recall that $\mathbf{e}_{1}$ is a random variable that’s the sum of $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-2^{\beta},2^{\beta}-1\right)}$ and $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$. By Lemma 7.1, there exists a positive absolute constant $c_{0}\in\mathbb{N}$ such that $\left\lVert\mathcal{D}\left(\mathbf{e}_{1}\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq c_{0}\cdot\frac{1}{\sqrt{\kappa}}$.
@@
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace,$
@@
  • The expectation of $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$ is $\mu_{i,0}=-\frac{1}{2}$ and the expectation of $\mathcal{U}^{(i)}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is $\mu_{i,1}=\frac{1}{2}$.
  • The variance of both $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$, $\mathcal{U}^{(i)}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is the same: $\sigma_{i,b}^{2}=\frac{\left(2^{\beta+1}+1\right)^{2}-1}{12}=\frac{2^{2\beta+2}+2^{\beta+2}}{12}=\Theta\left(2^{2\beta}\right)$.
@@
  • The third moment of both $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$, $\mathcal{U}^{(i)}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is of the same order:
    $\gamma_{i,0}:=\frac{\mathbb{E}_{u\leftarrow\mathcal{U}_{\left(-2^{\beta},2^{\beta}-1\right)}}\left(|u-\mu_{i,0}|^{3}\right)}{\sigma^{3}}=\Theta\left(\frac{1}{\kappa\sqrt{\kappa}}\right)\enspace,$
    $\gamma_{i,1}:=\frac{\mathbb{E}_{u\leftarrow\mathcal{U}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}}\left(|u-\mu_{i,1}|^{3}\right)}{\sigma^{3}}=\Theta\left(\frac{1}{\kappa\sqrt{\kappa}}\right)\enspace.$
@@
  • For every $i\in[\kappa/2],b\in\{0,1\}$, the random variable $S^{(i,b)}:=S-\mathcal{U}^{\left(i,b\right)}$ is summing all $\frac{\kappa}{2}$ variables from $S_{0}$ and $S_{1}$, except variable number $i$ from $S_{b}$.
@@
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)\right\rVert_{TV}$
$\leq\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\gamma_{i,b}+\frac{2\cdot\sigma_{i,b}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
@@
where $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$. We analyze each of the three summands separately, from last to first. As for the third (and last) summand, $\frac{1}{\sigma\cdot 2\sqrt{2\pi}}\leq O\left(\frac{1}{\sqrt{\kappa}\cdot 2^{\beta}}\right)$ because $\sigma=\Theta\left(\sqrt{\kappa}\cdot 2^{\beta}\right)$. The second summand is $\left(5+3\sqrt{\frac{\pi}{8}}\right)\cdot\gamma\leq O\left(\frac{1}{\sqrt{\kappa}}\right)$ because $\gamma=\Theta\left(\frac{1}{\sqrt{\kappa}}\right)$. Overall the sum of the third and second summands together is $O\left(\frac{1}{\sqrt{\kappa}}\right)$.
@@
$\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\gamma_{i,b}+\frac{2\cdot\sigma_{i,b}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
$=\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{i,b}+\frac{\sigma_{i,b}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
=$\underset{(*)}{=}\kappa\cdot\Theta\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{1,0}+\frac{\sigma_{1,0}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$=\Theta\left(\kappa\cdot\frac{3}{2}\cdot\sigma\cdot\gamma_{1,0}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$+\Theta\left(\kappa\cdot\frac{\sigma_{1,0}^{2}}{\sigma^{2}}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$\leq O\left(2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)+O\left(\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$\leq O\left(2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)\enspace,$
where the equality ()(*)( ∗ ) follows because for every $i\in[\kappa/2],b\in\{0,1\}$, $\gamma_{i,b}=\Theta\left(\gamma_{1,0}\right)$, $\sigma_{i,b}=\sigma_{1,0}$, and
$\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}=\Theta\left(\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)\enspace.$
Finally, by Fact 7.1,
$2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace,$
and overall we showed so far,
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace.$
@@
$\left\lVert\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}$
$\leq\left\lVert\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\tilde{B}\right)\right)-\mathcal{D}\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace.$
@@
Let $S$ and $S^{\left(i,b\right)}$ as defined in the statement and proof of Lemma 7.1. Then,
$\left\lVert\left(S^{\left(i,b\right)}\right)-\left(S^{\left(i,b\right)}+1\right)\right\rVert_{TV}\leq O\left(\frac{1}{2^{\beta}\cdot\sqrt{\kappa}}\right)\enspace.$
@@
$=e^{-\frac{\pi\cdot\left(\left\lVert\mathbf{e}_{0}\right\rVert^{2}+2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}\cdot B\right)}{2\cdot B^{2}}}\cdot\sum_{\mathbf{e}\in\mathbb{Z}^{n}_{\left(-B,\cdots,-1,0,1,\cdots,B\right)}}\frac{1}{P}\cdot e^{-\frac{\pi\cdot 2\cdot\left\lVert\mathbf{e}\right\rVert^{2}}{2\cdot B^{2}}}$
@@
Acknowledgements
We are grateful to Scott Aaronson and to Joseph Carolan for insightful discussions on the computational power of the models suggested in this work.